### v2

# A faithful action of Gal($\overline{\mathbb{Q}}/\mathbb{Q}$) on Zariski multiplets

Abstract.
In this work, we establish two main results in the context of arithmetic and geometric properties of plane curves. First, we construct numerous new examples of arithmetic Zariski pairs and multiplets, where only a few ones were previously available. Second, we describe a faithful action of the absolute Galois group on the equisingular strata of plane curves, providing insights into the interplay between Galois representations and the geometry of singular plane curves. We conclude the paper with very concretes examples of the general results obtained.
Key words and phrases:
Zariski pairs, absolute Galois group, moduli of surfaces of general type
2010 Mathematics Subject Classification:
14J10,14J29,20D15,20D25,20H10,30F99.
1. Introduction
The purpose of this work is twofold. First, we aim to expand the known landscape of arithmetic Zariski pairs and multiplets by providing a substantial number of new examples, enriching a domain where only a few ones were previously available [S08, AC17]. Second, we establish a faithful action of the absolute Galois group on the equisingular strata of plane curves, revealing an intriguing connection between arithmetic structures and singularity theory. Both results can be seen as a natural continuation of our previous work [LP20] adding the new results obtained in [BCG06, BCG15].
The notion of Zariski pair was introduced by E. Artal in his seminal paper [A94].
Definition 1.1.
A pair of complex reduced plane curves $(B_{1},B_{2})$ is called a Zariski pairs if it satisfies the following conditions:
  1. (1)
    There exist tubular neighborhoods $T(B_{i})$, $i=1,2$, and a homeomorphism $h\colon T(B_{1})\longrightarrow T(B_{2})$ such that $h(B_{1})=B_{2}$.
  2. (2)
    There exists no homeomorphism $f:\mathbb{P}^{2}\longrightarrow\mathbb{P}^{2}$ with $f(B_{1})=B_{2}$ i.e., the pairs $(\mathbb{P}^{2},B_{1})$ and $(\mathbb{P}^{2},B_{2})$ are not homeomorphic.
Analogously $(B_{1},\ldots,B_{k})$ is a Zariski $k$-plet if $(B_{i},B_{j})$ is a Zariski pair for any i $\neq j$.
As noted in [ACT10, Remark 3]) the first condition in Definition 1.1 can be replaced by equality of combinatorial data associated to the $B_{i}$, $i=1,2$, which in case of irreducible curves consist of the degree and the counting function associating to each topological type of singularities the number of corresponding singular points.
This definition put a new focus on an important question for plane complex projective curves:
Do the topological types of the singularities of a plane curve $B$ and the degree determine the homeomorphism type of the pair $(\mathbb{P}^{2},B)$?
Obviously this question has positive answer in case of plane curves without singularities, but already the case of nodal curves took a long time to be established beyond doubt, cf. Severi [S21], Harris [H86]. In these, and many other cases the proof shows the connectedness of the locus of curves with given singularity types in the projective space of plane curves of the given degree.
Instead, Zariski proved that in degree $6$ the complement of a sextic, regular except for six cusps, has non-abelian fundamental group only if the cusps lie on a conic, so to honour his results Zariski pairs were named.
The excellent survey [ACT10] can report on many new examples and provides some strategy to find additional ones in two steps
  1. (I)
    Locate curves of the same invariants in different connected components of the corresponding locus,
  2. (II)
    Find an effective invariant of the embedded pair, which distinguishes the curves.
In our previous work [LP20] we used the fundamental group of the complement together with the conjugacy class of meridians to the curve, which is well-defined for irreducible curves, in the second step.
Addressing the first step, the idea developed in [LP20] is to exploit the disconnectedness of moduli of surfaces of general type isogenous to a product. As these surface have ample canonical class with $K_{S}^{2}\geq 8$, the $m$-canonical map is an embedding into a projective space $\mathbb{P}^{N}$ at least for $m\geq 3$ by Bombieri’s theorem [B73].
An embedded surface then is mapped onto $\mathbb{P}^{2}$ using projection with center a disjoint projective subspace of codimension $3$. By the theorem of Ciliberto and Flamini [CiF11] the corresponding branch curve is reduced, irreducible and smooth except for ordinary nodes and cusps if the center is sufficiently general.
This procedure extends nicely to families of $m$-canonical embeddings and centers and thus associates a connected component of equisingular plane irreducible curves to a connected component of surfaces of general type with ample canonical bundle.
In this work we go even further. Inspired by and relying on the result given in [BCG15] we give new insight on arithmetic Zariski pairs. We define arithmetic Zariski pairs following Shimada
Definition 1.2.
A Zariski pair $(B_{1},B_{2})$ is called arithmetic Zariski pair if $B_{1}$ and $B_{2}$ are conjugate plane curves, i.e. there exists a polynomial $P(x_{0},x_{1},x_{2})$ with complex coefficients and an automorphism $\sigma$ of the field $\mathbb{C}$ such that
$B_{1}=\{P=0\},\textrm{ and }B_{2}=\{P^{\sigma}=0\}$
where $P^{\sigma}$ is the polynomial obtained from $P$ by applying $\sigma$ to its coefficients.
Our main theorem on Zariski multiplets of high cardinality follows our results [P13, GP14, LP14, LP16] on high cardinalities of the set of connected components of moduli spaces of surfaces isogenous to a product (SIP) and exploits the solution of the Chisini conjecture by Kulikov [K08] for generic branched projections.
Putting these into the perspective of the action of the absolute Galois group on varieties defined over number fields, we are going to show a few new results. The first concerns a faithful action of the absolute Galois group on loci of plane curves.
Theorem 1.3.
There is a faithful action of the absolute Galois group $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}$} on the set of connected components of equisingularity strata of plane irreducible curves with at most nodal and ordinary cuspidal singularities.
In particular we can derive from that set-up the existence of arbitrarily large arithmetic Zariski multiplets of plane curves. Indeed, we are able to make use of specific examples of large Galois orbits of surfaces for the final result.
Theorem 1.4.
There is an infinite sequence of prime numbers $p_{i}$ such that there is a Galois orbit of lengths $\lambda(p_{i})=\varphi(p^{2}_{i}-1)/8$ of branch curves $B_{i}$ of degree
$\deg(B_{i})=60p_{i}\left(\frac{p_{i}-1}{6}-1\right)\left(\frac{p_{i}+1}{4}-1\right).$
In particular the corresponding sequence of $\lambda(p_{i})$ is not bounded above.
Let us now explain the organization of this paper.
In the next two sections, addressing preliminary results, we lay the foundational tools and concepts necessary for our results. First, we define, for each automorphism $\sigma$ of the field of complex numbers $\mathbb{C}$, a functor $F_{\sigma}$ from the category of complex varieties to itself. This functor will serve as the central tool in our study of the action of the absolute Galois group on equisingular strata of plane curves. Second, we introduce the notion of generic $m$-canonical branch curves. These curves arise as branch curves of a covering $p\colon S\to\mathbb{P}^{2}$, where $S$ is a surface of general type. The covering $p$ is constructed as the composition of the $m$-canonical embedding of $S$ and a projection onto $\mathbb{P}^{2}$ from a generic disjoint center. We then recall the definition and basic properties of surfaces isogenous to a higher product and the weak rigidity theorem of Catanese [Cat00], which allows us to describe the connected components of the moduli space of such surfaces.
In the final section we first demonstrate that the associated generic $m$-canonical branch curves of surfaces isogenous to a product with different fundamental group form a Zariski pair (see Theorem 3.7) and, using, in particular cases, the functor $F_{\sigma}$, even arithmetic Zariski pairs. Finally, we provide detailed proofs of the main theorems stated in the introduction.
Acknowledgement The authors wish to thank F. Catanese, F. Flamini for fruitful discussions. The second author was partially supported by the Research Project PRIN 2020 - CuRVI, CUP J37G21000000001, the MIUR Excellence Department of Mathematics, University of Genoa, CUP D33C23001110001 and INDAM-GNSAGA.
2. Preliminaries
First, we investigate the action of ${\mathrm{Aut}}(\mathbb{C})$ and $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}$} on equisingular strata of the parametrizing space of plane curves with only nodes and cusps as singularities. Most of the material in this section is well known therefore we shall omit some details. For example a consequence of the Zorn Lemma is the following theorem.
Theorem 2.1.
Every $\sigma\in{\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}$ extends to an automorphism of the field of complex numbers $\mathbb{C}$. Hence we have a surjective morphism of groups
$\varphi\colon{\mathrm{Aut}}(\mathbb{C})\longrightarrow{\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}.$
Following [BCG15, Def.4.5] we consider functors on the category $\underline{\mathbf{\mathbb{C}-var}}$ of complex varieties.
Definition 2.2.
For each $\sigma\in Aut(\mathbb{C})$ let
$F_{\sigma}\colon\underline{\mathbf{\mathbb{C}-var}}$ $\longrightarrow$ $\underline{\mathbf{\mathbb{C}-var}}$
$X$ $\mapsto$ $X^{\sigma}$
$\left(f\colon X\longrightarrow Y\right)$ $\mapsto$ $\left(f^{\sigma}\colon X^{\sigma}\longrightarrow Y^{\sigma}\right)$
the functor that to a complex variety $X$ associates $X^{\sigma}=X\otimes_{\mathbb{C},\sigma}\mathbb{C}$.
The functor we have just defined shares the following properties (see [BCG15, section 4]):
Remark 2.3.
  1. (1)
    $\sigma\in{\mathrm{Aut}}(\mathbb{C})$ acts on $\mathbb{C}[z_{0},\dots,z_{n}]$ by sending the element $P(z)=\sum_{I=(i_{0},\dots,i_{n})}a_{I}z^{I}$ to
    $\sigma(P)(z)\quad:=\quad\sum_{I=(i_{0},\dots,i_{n})}\sigma(a_{I})z^{I}.$
  2. (2)
    Let $X$ be a projective variety
    $X\quad\subset\quad\mathbb{P}^{n}_{\mathbb{C}},\quad X:=\{z\mid f_{i}(z)=0\;\forall i\,\}.$
    The action of $\sigma$ extends coordinatewise to $\mathbb{P}^{n}_{\mathbb{C}}$ and carries $X$ to the set $\sigma(X)$ which is another variety that coincides with $X^{\sigma}$, the conjugate variety.
    In fact, since $f_{i}(z)=0$ if and only if $\sigma(f_{i})(\sigma(z))=0$, one has that
    $X^{\sigma}\quad=\quad\{w\mid\sigma(f_{i})(w)=0\;\forall i\,\}.$
  3. (3)
    If $X$ is defined over a subfield $k_{0}\subset\mathbb{C}$, i.e., there is a $k_{0}$-scheme $X_{0}$ such that $X\cong X_{0}\otimes_{k_{0}}\mathbb{C}$, then $X^{\sigma}$ depends only on the restriction $\sigma|_{k_{0}}$ and if moreover $\sigma$ is the identity on $k_{0}$, then $X^{\sigma}$ is canonically isomorphic to $X$; we will use this quite often with $k_{0}=\mathbb{Q}$.
  4. (4)
    The formation of $X^{\sigma}$ is compatible with products;
  5. (5)
    For a group action $G\times X\to X$ with $G$ finite (hence defined over $\mathbb{Q}$ as an algebraic group), then $G^{\sigma}=G$ canonically and applying the conjugation functor gives a conjugate action $G\times X^{\sigma}=G^{\sigma}\times X^{\sigma}\to X^{\sigma}$.
In [BCG15, section 5] the authors argue that the action of ${\mathrm{Aut}}(\mathbb{C})$ on all complex surfaces descents to an action of ${\mathrm{Aut}}(\mathbb{C})$ on the concected components of the moduli space of complex projective surfaces of general type. Indeed if $S_{1}$ and $S_{2}$ are two surfaces belonging to the same family $\mathcal{X}\longrightarrow B$ we can apply the functor $F_{\sigma}$ and get the following functorial diagram
[…]
and the family $\mathcal{X}^{\sigma}\longrightarrow B^{\sigma}$ connects the surfaces $S_{1}^{\sigma}$ and $S_{2}^{\sigma}$. More precisely, choose the canonical model $X$ of a surface of general type $S$ and apply the field automorphism $\sigma\in{\mathrm{Aut}}(\mathbb{C})$ to a point of the Hilbert scheme corresponding to the $m$-canonical image of $S$ . We obtain a surface $X^{\sigma}$, and whose minimal model is $S^{\sigma}$. Since the Hilbert scheme corresponding to $m$-pluricanonical embedded surfaces is defined over $\mathbb{Z}$, it is invariant under the functor $F_{\sigma}$ so its connected components are defined over $\overline{\mathbb{Q}}$. We conclude that the elements in the kernel of $\varphi\colon{\mathrm{Aut}}(\mathbb{C})\longrightarrow{\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}$ act trivially on connected components of the moduli space of surfaces which is in bijection with the set of connected components of the Hilbert subscheme just defined. For more details see [BCG15, Proposition 5.7].
We want to do something similar for parameter spaces of plane curves. More precisely, to replace the Hilbert scheme we consider equisingularity strata of plane curves with a fixed combinatorial type. Like the Hilbert scheme, such strata are in general neither irreducible nor connected.
Definition 2.4.
Given a combinatorial type of degree $d$ plane curves as defined in full generality in [ACT10], the corresponding equisingular stratum in the parameter space of degree $d$ plane curve is given by the locus of all curves with that combinatorial type.
Below we will only consider irreducible plane curves with ordinary nodes and cusps with topological type determined by the respective Milnor numbers $\mu=1$ and $\mu=2$. In this situation, the combinatorial datum of $B$ is simply encoded by the triple $(d,n,c)$, where $d=\deg B$ is the degree, $n$ is the number of nodes, and $c$ is the number of cusps of $B$.
The corresponding equisingular stratum is a quasi-projective variety
$\mathcal{S}^{d}_{n,c}\subset\mathbb{P}^{\frac{d(d+3)}{2}}=:\mathbb{P},$
as proved by Wahl in [W74]. In particular, $\mathcal{S}^{d}_{n,c}$ is a locally closed subset of a projective space, and hence a complex algebraic variety endowed with the induced standard analytic topology. Since such spaces are locally path connected, the connected components of $\mathcal{S}^{d}_{n,c}$ coincide with its path-connected components. Consequently, $\mathcal{S}^{d}_{n,c}$ has finitely many (path-)connected components.
Explicitly, we can describe
[…]
We observe that if $(B_{1},B_{2})$ is a Zariski pair, where each $B_{i}$ is given by an equation $\{p_{i}=0\}$, $i=1,2$, then the corresponding points $p_{1}]$ and $p_{2}]$ lie in different path-connected components of $\mathcal{S}^{d}_{n,c}$. The converse is not known in general.
This motivates the following definition where, by abuse of notation, we shall consider a curve $B_{i}$ also as element in $\mathcal{S}^{d}_{n,c}$.
Definition 2.5.
Two curves $B_{1},B_{2}$ in $\mathcal{S}^{d}_{n,c}$ are said to be rigidly isotopic if there exists a continuous path (in the standard analytic topology) $\gamma:[0,1]\to\mathcal{S}^{d}_{n,c}$ such that $\gamma(0)=[p_{1}]$ and $\gamma(1)=[p_{2}]$, where $B_{i}=\{p_{i}=0\}$ for $i=1,2$. Equivalently, $B_{1}$ and $B_{2}$ belong to the same path-connected component of the equisingular stratum $\mathcal{S}^{d}_{n,c}$.
Remark 2.6.
To get this definition into the right perspective let us note the following:
  1. (1)
    In [ACT10], rigid isotopy is defined using smooth paths. We observe that, in this context, the two definitions – using either continuous paths or smooth paths – define the same path components, hence the same equivalence relation.
  2. (2)
    From an algebraic viewpoint, such a path corresponds to a flat projective family of plane curves $\pi:\mathcal{X}\to T$ over a connected complex curve $T$, whose fibers $\mathcal{X}_{t}\subset\mathbb{P}^{2}$ all have degree $d$ and precisely $n$ nodes and $c$ cusps. Thus, rigid isotopy classes coincide with the connected components of the equisingular stratum, both topologically and algebraically.
  3. (3)
    Consequently, the connected components of the equisingular stratum $\mathcal{S}^{d}_{n,c}$ are precisely the rigid isotopy classes of curves.
We now unpack the last two remarks more carefully, since they involve both the analytic and the Zariski topologies.
Let $p_{1}],[p_{2}]\in\mathcal{S}^{d}_{n,c}\subset\mathbb{P}^{\frac{d(d+3)}{2}}$. We define the following equivalence relations reflecting the two topologies:
  1. (1)
    $p_{1}]\sim_{1}[p_{2}]$ if they lie in the same path-connected component of $\mathcal{S}^{d}_{n,c}$ in the standard analytic topology;
  2. (2)
    $p_{1}]\sim_{2}[p_{2}]$ if there exists a finite chain of quasi-projective irreducible curves $T_{1},\dots,T_{m}\subset\mathcal{S}^{d}_{n,c}$ such that $T_{i}\cap T_{i+1}\neq\varnothing$ for all $i$ and $p_{1}],[p_{2}]$ are points on $T_{1}$ resp. $T_{m}$.
    (Equivalently, $\sim_{2}$ is the transitive closure of the relation “lies on a common irreducible quasi-projective curve contained in $\mathcal{S}^{d}_{n,c}$”).
Then these two equivalence relations coincide:
$p_{1}]\sim_{1}[p_{2}]\quad\Longleftrightarrow\quad[p_{1}]\sim_{2}[p_{2}].$
Indeed, the implication from $\sim_{2}$ to $\sim_{1}$ is clear, since each irreducible algebraic curve $T_{i}$ is path connected with the analytic topology.
For the converse, assume $p_{1}]\sim_{1}[p_{2}]$ and let $\gamma:[0,1]\to\mathcal{S}^{d}_{n,c}$ be a path joining them, continuous in the analytic topology of $\mathcal{S}^{d}_{n,c}$. Denote by $S_{1},\dots,S_{\ell}$ the irreducible components of $\mathcal{S}^{d}_{n,c}$ with respect to the Zariski topology. Then for every path $\gamma$ as above, there exist a sequence $i_{1},\dots,i_{k}$ of elements in $\{1,\dots,\ell\}$ and real numbers $0=t_{0}<t_{1}<\dots<t_{k}=1$ such that $\gamma([t_{j-1},t_{j}])\subset S_{i_{j}}$.
By induction on $k\geq 1$ the claim then follows: If $k=1$, both $p_{1}],[p_{2}]$ belong to the same $S_{i_{1}}$ and an irreducible algebraic curve exists on $S_{i_{1}}$ which contains both.
If $k>1$ then by induction hypothesis both $p_{1}],\gamma(t_{k-1})$ belong to a connected union of algebraic curves. Since $\gamma(t_{k-1}),[p_{2}]$ both belong to the same component $S_{i_{k}}$ the inductive claim also follows, thus $p_{1}]\sim_{2}[p_{2}]$.
Moreover, if $p_{1}]\sim_{2}[p_{2}]$ via a connected quasi-projective curve $T\subset\mathcal{S}^{d}_{n,c}$, we can construct the corresponding flat family as follows: consider the universal family $\mathcal{C}\subset\mathbb{P}^{2}\times\mathbb{P}^{\frac{d(d+3)}{2}}$ of irreducible degree-$d$ curves, and let
(2) $\pi:\mathcal{X}:=\mathcal{C}\times_{\mathbb{P}^{\frac{d(d+3)}{2}}}T\longrightarrow T$
be the induced family. Then $\pi$ is a flat projective family of plane curves of degree $d$ with exactly $n$ nodes and $c$ cusps as singularities.
In a similar way as for surfaces we use the diagram (1) in this context, in particular we refer to the map (2) and get:
Proposition 2.7.
For each $\sigma\in{\mathrm{Aut}}(\mathbb{C})$ the functor $F_{\sigma}$ descends to a well defined map on rigid isotopic classes of plane curves.
By this proposition we can define an action of ${\mathrm{Aut}}(\mathbb{C})$ on the set of all rigid isotopic classes of plane curves.
A consequence of the Theorem [PR20, Theorem 11] of Parusinski–Rond is that each such class contains a curve defined over $\bar{\mathbb{Q}}$, and we get the following result
Proposition 2.8.
For each $\sigma$ in the kernel of ${\mathrm{Aut}}(\mathbb{C})\to{\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}$ the functor $F_{\sigma}$ preserves all rigid isotopy classes of irreducible plane curves.
Hence $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}$} acts on the rigid isotopy classes.
Proof.
The image under $F_{\sigma}$ of the rigid isotopy class of a curve $B$ is the isotopy class of the image of any curve rigidly isotopic to $B$ under $F_{\sigma}$. Since $\sigma$ is in the kernel of ${\mathrm{Aut}}(\mathbb{C})\to{\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}$, this functor fixes all curves defined over $\bar{\mathbb{Q}}$. Now every isotopy class contains such curve by [PR20, Theorem 11 ii) & iii)], $F_{\sigma}$, hence $\sigma$ fixes every isotopy class. Thus the ${\mathrm{Aut}}(\mathbb{C})$-action descends to a $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}$}-action on isotopy classes. ∎
We want to link more tightly the action of $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}$} on the connected components of moduli space of surfaces with the action on the connected components of equisingular strata of plane curves with only nodes and cusps as singularities. To do so we have to recall the general theory of branched curves, this is the same strategy we used in our previous paper [LP20].
3. branch curves and surfaces isogenous to a product
There is natural way to produce many singular plane curves with the above mentioned singularities. Indeed, it is enough to consider branch curves of coverings of $\mathbb{P}^{2}$. To give such covering we will proceed in the following way: first we consider a surface of general type $S$ with ample canonical class, then we consider the natural immersion in a $\mathbb{P}^{n}$ by a multicanonical system. Finally, we project the image generically to $\mathbb{P}^{2}$. This yields a covering $S\longrightarrow\mathbb{P}^{2}$. To be more precise let us explain in details this procedure.
Definition 3.1 (cf. [K99]).
Let $B\subset\mathbb{P}^{2}$ be an irreducible plane algebraic curve with ordinary cusps and nodes as the only singularities. The curve $B$ is called generic branch curve if there is a finite morphism $p:S\to\mathbb{P}^{2}$ with $\deg p\geq 3$ such that
  1. (1)
    $S$ is a smooth irreducible projective surface,
  2. (2)
    $p$ is unramified over $\mathbb{P}\setminus B$,
  3. (3)
    $p^{*}(B)=2R+C$, where $R$ is a smooth irreducible reduced curve and $C$ is a reduced curve,
  4. (4)
    the morphism $p_{|R}:R\to B$ coincides with the normalization of $B$.
In this case $p$ is called a generic covering of the projective plane.
Theorem 3.2.
[CiF11, Theorem 1.1] Let $S\subset\mathbb{P}^{r}$ be a smooth, irreducible, projective surface. Then the ramification curve on $S$ of a generic projection of $S$ to $\mathbb{P}^{2}$ is smooth and irreducible and the branch curve in the plane is also irreducible and has only nodes and cusps, respectively, corresponding to two simple ramification points and one double ramification point.
In particular such generic projection of $S$ onto $\mathbb{P}^{2}$ has a generic branch curve in the sense of Def. 3.1.
Definition 3.3.
A plane curve $B$ is called a generic m-canonical branch curve if there exists a smooth surface $S$ with $mK_{S}$ very ample and a commutative diagram
[…]
where $\phi_{m}$ is the $m$-canonical map, $p$ is a generic covering and $B$ is the branch locus of $p$.
We are now able to associate to a surface of general type a plane curve with only nodes and cusps as singularities. Among all the surfaces of general type there are some for which it is possible to control accurately the $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}$} action on the connected components of their moduli space. These surfaces are the so called surfaces isogenous to a product (or SIPs for short). They were introduced by Catanese in [Cat00] and the action of $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}$} on their moduli space was investigated in [BCG15].
Thanks to their simple definition SIPs are incredibly versatile and they give rise to a large amount of interesting examples. Let us be more precise:
Definition 3.4.
A surface $S$ is said to be isogenous to a higher product of curves if and only if, $S$ is a quotient $(C_{1}\times C_{2})/G$, where $C_{1}$ and $C_{2}$ are curves of genus at least two, and $G$ is a finite group acting freely on $C_{1}\times C_{2}$. (To ease notation we call the surfaces SIP)
Using the same notation as in Definition 3.4, let $S$ be a SIP, and $G^{\circ}:=G\cap(Aut(C_{1})\times Aut(C_{2}))$. Then $G^{\circ}$ acts on the two factors $C_{1}$, $C_{2}$ and diagonally on the product $C_{1}\times C_{2}$. If $G^{\circ}$ acts faithfully on both curves, we say that $S=(C_{1}\times C_{2})/G$ is a minimal realization. In [Cat00] it is also proven that any SIP admits a unique minimal realization.
Assumptions. In the following we always assume:
  1. (1)
    Any SIP $S$ is given by its unique minimal realization;
  2. (2)
    $G^{\circ}=G$, this case is also known as unmixed type, see [Cat00].
Under these assumptions we have.
Proposition 3.5.
[Cat00] Let $S=(C_{1}\times C_{2})/G$ be a SIP, then $S$ is a minimal surface of general type with the following invariants:
(3) $\chi(S)=\frac{(g(C_{1})-1)(g(C_{2})-1)}{|G|},\quad e(S)=4\chi(S),\quad K^{2}_{S}=8\chi(S).$
The irregularity of these surfaces is computed by
(4) $q(S)=g(C_{1}/G)+g(C_{2}/G).$
Moreover the fundamental group $\pi_{1}(S)$ fits in the following short exact sequence of groups
(5) $1\longrightarrow\pi_{1}(C_{1})\times\pi_{1}(C_{2})\longrightarrow\pi_{1}(S)\longrightarrow G\longrightarrow 1.$
The most important property of surfaces isogenous to a product is their weak rigidity property.
Theorem 3.6.
[Cat03, Theorem 3.3, Weak Rigidity Theorem] Let $S=(C_{1}\times C_{2})/G$ be a surface isogenous to a higher product of curves. Then every surface with the same
  • topological Euler number and
  • fundamental group
is diffeomorphic to $S$. The corresponding moduli space $\mathcal{M}^{top}(S)=\mathcal{M}^{\it diff}(S)$ of surfaces (orientedly) homeomorphic (resp. diffeomorphic) to $S$ is either irreducible and connected or consists of two irreducible connected components exchanged by complex conjugation.
We can now link SIP and branch curves as done in [LP20]. In this contest we need to be more careful in choosing the pluricanonical embeddings because we have no specific restrictions on the group $G$ as we had in [LP20]. To this end we have change the hypothesis of [LP20, Lemma 3.6] and the claim is proved for the $3$-canonical embedding instead of $2$-canonical. We get an analogue theorem to [LP20, Theorem 3.12].
Theorem 3.7.
Let $S_{1}$ and $S_{2}$ be two surfaces isogenous to a product with the same Euler number and $\pi_{1}(S_{1})\neq\pi_{1}(S_{2})$. Then for $m\geq 3$ the corresponding generic m-canonical branch curves $(\mathbb{P}^{2},B_{1})$, $(\mathbb{P}^{2},B_{2})$ are a Zariski pair.
Now that we have a link between Zariski pairs and SIPs we want to add arithmetic information to this setting. In [BCG15] the three authors exploited the weak rigidity theorem to prove the following result.
Theorem 3.8.
[BCG15, Theorem 1.1] If $\sigma\in{\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}$ is not in the conjugacy class of the complex conjugation, then there exists a surface isogenous to a product $X$ such that $X$ and the Galois conjugate varieties $X^{\sigma}$ have non-isomorphic fundamental group.
Moreover, they could prove even a much stronger theorem.
Theorem 3.9.
[BCG15, Theorem 1.3] The absolute Galois group $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}$} acts faithfully on the set of connected components of the (coarse) moduli space of surfaces of general type.
We conclude this preliminary section with a consideration of numerical invariants. Since these numerical invariants of a SIP are completely determined by the Euler number, which is a fortiori determined by the genera of the curves involved and by the order of the group $G$, we are able to determine the degree $d$, the number $n$ of nodes, and the number $c$ of cusps of the branch curve. Indeed repeating the same argument as for the proof of [LP20, Lemma 3.10] we have:
Lemma 3.10.
Let $(\mathbb{P}^{2},B)$ be a generic 3-canonical branch curve of regular surfaces isogenous to a product $S$ then
(6) $d=\deg B=30c_{1}(S)^{2},$
(7) $c=137c_{1}(S)^{2}-c_{2}(S)$
(8) $n=450(c_{1}(S)^{2})^{2}-237c_{1}(S)^{2}+c_{2}(S).$
4. Proof of the main theorems
To prepare the proofs we combine the results in Theorem 3.7 and Theorem 3.8 to give examples of arithmetic Zariski pairs as follows.
Theorem 4.1.
Let $S$ and $S^{\sigma}$ be two surfaces isogenous to a product as in Theorem 3.8 then they have the same Euler number but $\pi_{1}(S)\neq\pi_{1}(S^{\sigma})$. Then for $m\geq 3$ the corresponding generic m-canonical branch curves $(\mathbb{P}^{2},B)$, $(\mathbb{P}^{2},B^{\sigma})$ are arithmetic Zariski pair.
Proof.
By Theorem 3.7 $(\mathbb{P}^{2},B)$, $(\mathbb{P}^{2},B^{\sigma})$ are Zariski pair. We only have to prove that the pair is arithmetic. By the proof of [BCG15, Theorem 1.3] the field of moduli is a number field. Also a field of definition is a number field, though a bigger one in general, therefore we know that the $m$-canonical model $X$ of $S$ and $X^{\sigma}$ for $S^{\sigma}$ in $\mathbb{P}^{P_{m}-1}$ is defined over $\overline{\mathbb{Q}}$. By [CiF11, Theorem 1.1] we can map $S$ to $\mathbb{P}^{2}$ by a generic projection $\pi$ with branch curve $B$ that is irreducible with at most nodes and ordinary cusps as singularities and again defined over a number field. In particular, this allows us to apply the functor $F_{\sigma}$ defined in 2.2 and to obtain the following diagram.
${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${X^{\sigma}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{\sigma}}$$\scriptstyle{F_{\sigma}}$${B\subset\mathbb{P}^{2}}$${B^{\sigma}\subset\mathbb{P}^{2}}$
This completes the proof, since also $\pi^{\sigma}$ is a generic projection and $B^{\sigma}$ is a generic branch curve. ∎
We are now in the position to prove Theorems 1.3 and 1.4, that we recall for reader’s convenience.
Theorem 4.2.
There is a faithful action of the absolute Galois group $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}$} on the set of connected components of equisingularity strata of plane irreducible curves with at most nodal and ordinary cuspidal singularities.
Proof.
For every $\sigma\in{\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}$ non trivial, there exists a SIP $X$ defined over $\overline{\mathbb{Q}}$ such that $X$ and $X^{\sigma}$ have non isomorphic topological fundamental group. This follows from Theorem 3.8.
We can embed $X$ by the $3$-canonical map into a projective space then consider a linear projection $\pi$ from a generic center of projection to $\mathbb{P}^{2}$. This gives a branch curve $B$, that is irreducible with at most nodes and ordinary cusps as singularities by Theorem 3.2.
We extend $\sigma$ to an automorphism of $\mathbb{C}$ to get another projection $\pi^{\sigma}\colon X^{\sigma}\longrightarrow\mathbb{P}^{2}$ branched along $B^{\sigma}$. Then by our previous result [LP20] $(\mathbb{P}^{2},B)$ and $(\mathbb{P}^{2},B^{\sigma})$ is a Zariski pair by Theorem 4.1 . We deduce that $B$ and $B^{\sigma}$ are in different strata and thus $\sigma$ does not act trivially on the set of strata. ∎
Finally we use the article [GJT18] to give more concrete examples of larger multiplets. Their construction involves Beauville surfaces, which are rigid SIP and their result is stated in terms of the Euler totient function $\varphi$:
Theorem 4.3.
There is an infinite sequence of prime numbers $p_{i}$ such that there is a Galois orbit of lengths $\lambda(p_{i})=\varphi(p^{2}_{i}-1)/8$ of branch curves $B_{i}$ of degree
$\deg(B_{i})=60p_{i}\left(\frac{p_{i}-1}{6}-1\right)\left(\frac{p_{i}+1}{4}-1\right).$
In particular the corresponding sequence of $\lambda(p_{i})$ is not bounded above.
Proof.
By Dirichlet the sequence of prime numbers $p$ such that $p\equiv 19\,(\mod 24)$ is not bounded. Since $\varphi$ takes every integer value at most a finite number of times the last claim is immediate.
For each one of these primes $p$, we know by Theorem 7.1 [GJT18] there exists a $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}$} orbit of Beauville surfaces with branching data
$(2,3,p-1),\quad(2,4,p+1)\quad\textrm{ and group}\quad{\mathrm{PGL}}(2,p)$
of length
$\lambda(p)=\frac{1}{8}\cdot\varphi\left(p^{2}-1\right).$
We use the formula given in Lemma 3.10 to get that the degrees of the corresponding branch curves $B_{p}$ are
$\deg(B_{p})=60p\left(\frac{p-1}{6}-1\right)\left(\frac{p+1}{4}-1\right).$
To get a Galois orbit of the same length we have to show that the branch curves of two surfaces in different components are mapped to different components of strata.
Assume the contrary that they belong to the same stratum then by our result [LP20, Theorem 3.12] also the surfaces have isomorphic fundamental group, if so by the weak rigidity theorem they are in the same component contradicting our assumption. ∎